OpenStudy (anonymous):

solve by using system of equations 6x - 10y =8 2x + 4y =7

OpenStudy (callielovesyhuu):

y=7/4+1/2x

OpenStudy (jhannybean):

For this I think the elimination method would work best. You've got 2 equations: 6x - 10 = 8, 2x + 4 = 7 We can eliminate one variable, and solve for the other. 6x - 10 = 8 2x + 4 = 7 First, we decide which variable would be easier to eliminate. In this case, since 6 is a multiple of 2, we can use this to our advantage. We know that in order to eliminate the 6x we need to change 2x to -6x.

OpenStudy (jhannybean):

Sorry, i forgot to add in my y's. 6x - 10y = 8 2x + 4y = 7 *

OpenStudy (jhannybean):

How would we do this? We multiply the entire second equation by -3. -3(2x + 4y = 7) this gives us -6x -12y = -21 Using this derived equation, we can eliminate the x's. 6x - 10y = 8 -6x -12y = -21 --------------- 0 - (10+12)y = 8-21 With this we have our equation for y. -22y = -13 y = -13/ -21 y = 13/21

OpenStudy (jhannybean):

Now we use our value of y and substitute it into one of the given equations to find x. Let's take the second one for example. When y = 13/21, 2x + 4(13/21) = 7 distribute 2x + (4 * 13)/21 = 7 2x + 52/21 = 7 keep all the x's on one side and all your constants on the other. 2x = 7 - (52/21) 2x = 95/21 x = (95/21)/2 x = 95/42